# BeeManc at the PLABA Track of TAC-2023: Investigating LLMs and Controllable Attributes for Improving Biomedical Text Readability

### v1

@@ Abstract
In this system report, we describe the models and methods we used for our participation in the PLABA2023 task on biomedical abstract simplification, part of the TAC 2023 tracks. The system outputs we submitted come from the following three categories: 1) domain fine-tuned T5-like models including Biomedical-T5 and Lay-SciFive; 2) fine-tuned BART-Large model with controllable attributes (via tokens) BART-w-CTs; 3) ChatGPT-prompting. We also present the work we carried out for this task on BioGPT fine-tuning.
@@
In the official automatic evaluation using SARI scores, BeeManc ranks 2nd among all teams and our model Lay-SciFive ranks 3rd among all 13 evaluated systems. In the official human evaluation, our model BART-w-CTs ranks 2nd on Sentence-Simplicity (score 92.84), 3rd on Term-Simplicity (score 82.33) among all 7 evaluated systems; It also produced a high score 91.57 on Fluency in comparison to the highest score 93.53. In the second round of submissions, our team using ChatGPT-prompting ranks the 2nd in several categories including simplified term accuracy score 92.26 and completeness score 96.58, and a very similar score on faithfulness score 95.3 to re-evaluated PLABA-base-1 (95.73) via human evaluations.
Our codes, fine-tuned models, prompts, and data splits from the system development stage will be available at https://github.com/HECTA-UoM/PLABA-MU
1 Introducing the Methodologies
@@
The overall framework of our experimental design is displayed in Figure 1. In the first step, we fine-tune selected Large Language Models (LLMs) including T5, SciFive, BioGPT, and BART, apply prompt-based learning for ChatGPTs, and optimise control mechanisms on BART model. Then, we select the best performing two models using quantitative evaluation metrics SARI, BERTscore, BLEU, and ROUGE. Finally, we chose a subset of the testing results of the two best-performing models for our human evaluation. In this section, we first introduce the models we used, followed by LoRA efficient training, and then introduce the quantitative metrics we applied.
@@
Figure 1: Model Development and Selection. BART is fine-tuned using Wikilarge data. MAX step chooses the two best-performing models according to the automatic evaluation results using SARI and BERTscore.
1.1 Models
@@
The baseline models we investigated in our work include T5, SciFive, GPTs, BioGPT, BART, and Control Mechanisms. We give more details below.
1.1.1 T5
@@
T5 Raffel et al. (2020) used the same Transformer structure from Vaswani et al. (2017) but framed the text-to-text learning tasks using the same vocabulary, sentence piece tokenisation, training, loss, and decoding. The pre-fixed tasks include summarisation, question answering, classification, and translation. T5 used the common crawl corpus and filtered to keep only natural text and de-duplication processing. They extracted 750GB of clean English data to feed into the model for multi-task pre-training. Different masking strategies are integrated into the T5 model to facilitate better performances of specific fine-tuning tasks. Notably, it has been successfully employed in various fields such as Clinical T5 by Lehman and Johnson (2023) and Lu et al. (2022).
@@
Furthermore, T5’s pre-training on an extensive and diverse corpus of text data endows it with a strong foundation in understanding intricate language structures, a crucial asset for handling medical terminology. Its fine-tuning capability further enhances its adaptability, allowing us to tailor the model specifically to the nuances of the biomedical text simplification task. These attributes make T5 a compelling candidate for this work. In this paper, we fine-tuned three versions of T5, namely t5-small, t5-base, and t5-large, paired with their sentence-piece pre-trained tokenizer, which leads to our Biomedical-T5s. Each is fine-tuned independently on the same dataset as the other models to provide comparable results. Note that we use the prompt “summarize:” as it is the closest to our task.
1.1.2 SciFive
@@
Using the framework of T5, SciFive is a Large Language Model pre-trained on the biomedical domain and has demonstrated advanced performances on multiple biomedical NLP tasks Phan et al. (2021).
While preserving the abilities of T5 in sequence-to-sequence tasks, SciFive offers a deep understanding of medical terminology, concepts, and language structures. As a result, SciFive emerges as a strong candidate for text summarisation and medical language processing tasks, offering the potential to generate clear and accurate simplifications of medical texts.
@@
Similarly to our work on T5, we fine-tuned two versions of SciFive, namely SciFive-base and SciFive-large, paired with their pre-trained tokenizer. Each is fine-tuned independently on the same dataset. We again use the prompt “summarize:” for task-relation purposes. Because SciFive itself was already pre-trained in the biomedical domain, we name our fine-tuned SciFive for text simplification task as “Lay-SciFive”.
Figure 2: GPT Prompt Examples
1.1.3 OpenAI’s GPTs
@@
Given the remarkable performance demonstrated by OpenAI’s GPT models in text simplificationJeblick et al. (2022), we decided to apply simplifications using “GPT-3.5-turbo” and GPT-4 via their API [^1] https://openai.com/blog/openai-api. Both models were accessed. Example prompts we used can be found in Figure 2.
1.1.4 BioGPT
@@
BioGPT Luo et al. (2022) is an advanced language model specifically designed for medical text generation. BioGPT is built upon GPT-3 but is specifically trained to understand medical language, terminology, and concepts. BioGPT follows the Transformer language model backbone and is pre-trained on 15 million PubMed abstracts. It has demonstrated a high level of accuracy and has great potential for applications in medicine.
BioGPT is fine-tuned on our training and validation set, as with other encoder-decoder models (Figure 1) producing our Lay-BioGPT.
1.1.5 BART
@@
Like the default transformer structure, BART Lewis et al. (2020) aims to address the issues in BERT and GPT models by integrating their structures with a bi-directional encoder and an autoregressive decoder. In addition to the single token masking strategy applied in BERT, BART provides various masking strategies, including deletion, span masking, permutation, and rotation of sentences. Compared to GPT models, BART provides both leftward and rightward context in the encoders.
1.1.6 Controllable Mechanisms
@@
We applied the modified control token strategy in Li et al. (2022) for both BART-base and BART-large models. The training includes 2 stages, leveraging both Wikilarge training set Zhang and Lapata (2017) and our split of training set from PLABA Attal et al. (2023).
@@
They represent 1) the syntactic complexity, 2) the lexical complexity, 3) the inverse similarity of input and output at the character level, and 4) the length ratio of input and output respectively.
@@
After fine-tuning, the next step is to find the optimal value of CTs. Following a similar process in MUSS Martin et al. (2020), we applied Nevergrad Rapin and Teytaud (2018) on the validation set to find the static optimal discrete value for DTD, WR, and LV. As for the LR, we applied the control token predictor to maximise the performance with the flexible value. The predictor is also trained on Wikilarge data Zhang and Lapata (2017) to predict the potential optimal value for LR. We call the integrated methods of applying CTs on the BART model as BART-w-CTs.
1.2 LoRA and LLMs
@@
To evaluate bigger model architectures, we fine-tune FLAN-T5 XL Chung et al. (2022) and BioGPT-Large, which have 3 billion and 1.5 billion parameters, respectively. FLAN-T5 XL is based on the pre-trained T5 model with instructions for better zero-shot and few-shot performance. To optimise training efficiency, and as our computational resources do not allow us to fine-tune the full version of these models, we employ the LoRA Hu et al. (2022) technique, which allows us to freeze certain parameters, resulting in more efficient fine-tuning with minimal trade-offs.
2 Model Development
The only data set we used for the model fine-tuning development is from PLABA Attal et al. (2023). So, our team belongs to the data-constrained category if other teams applied broader and external biomedical data for their model development.
2.1 Data Preprocessing and Setup
@@
To investigate the selected models for training and fine-tuning, we divided the original PLABA data into Train, Validation, and Test sets, aiming for an 8:1:1 ratio. However, in the real implementations, we found that there are only a few 1-to-0 sentence pairs, which might cause a negative effect in training the simplification models. Thus we eliminated all 1-to-0 sentence pairs. In addition, to better leverage the SARI score, we picked sentences with multi-references for validation and testing purposes. As a result, we ended up with the following sentence pair numbers according to the source sentences (5757, 814, 814). Note that for our submitted three system outputs from BART-w-CTs, Biomedical-T5, and Lay-SciFive, the first one only used our training set split, while the other two used both our training and validation sets. We expect the models will have even better performances if we have re-trained the model using all the PLABA data with the optimised parameters, i.e. including our test set split.
2.2 Automatic Evaluations

### v2

# Large Language Models for Biomedical Text Simplification: Promising But Not There Yet

@@ Abstract
Biomedical literature often uses complex language and inaccessible professional terminologies. That is why simplification plays an important role in improving public health literacy. Applying Natural Language Processing (NLP) models to automate such tasks allows for quick and direct accessibility for lay readers. In this work, we investigate the ability of state-of-the-art large language models (LLMs) on the task of biomedical abstract simplification, using the publicly available dataset for plain language adaptation of biomedical abstracts (PLABA). The methods applied include domain fine-tuning and prompt-based learning (PBL) on: 1) Encoder-decoder models (T5, SciFive, and BART), 2) Decoder-only GPT models (GPT-3.5 and GPT-4) from OpenAI and BioGPT, and 3) Control-token mechanisms on BART-based models. We used a range of automatic evaluation metrics, including BLEU, ROUGE, SARI, and BERTScore, and also conducted human evaluations. BART-Large with Control Token (BART-L-w-CT) mechanisms reported the highest SARI score of 46.54 and T5-base reported the highest BERTScore 72.62. In our internal human evaluation, BART-L-w-CTs achieved a better simplicity score over T5-Base (2.9 vs. 2.2), while T5-Base achieved a better meaning preservation score over BART-L-w-CTs (3.1 vs. 2.6).
@@
We report the submission to PLABA2023 shared task. In the official automatic evaluation using SARI scores, BeeManc ranks 2nd among all teams and our model Lay-SciFive ranks 3rd among all 13 evaluated systems. In the official human evaluation, our model BART-w-CTs ranks 2nd on Sentence-Simplicity (score 92.84), 3rd on Term-Simplicity (score 82.33) among all 7 evaluated systems; It also produced a high score 91.57 on Fluency in comparison to the highest score 93.53. In the second round of submissions, our team using ChatGPT-prompting ranks the 2nd in several categories including simplified term accuracy score 92.26 and completeness score 96.58, and a very similar score on faithfulness score 95.3 to re-evaluated PLABA-base-1 (95.73) via human evaluations. Our codes, fine-tuned models, and data splits from the system development stage will be available at https://github.com/HECTA-UoM/PLABA-MU
Index Terms:
Large Language Models, Text Simplification, Biomedical NLP, Control Mechanisms, Health Informatics
I Introduction
The World Health Organization (WHO) defines health literacy as: “the personal characteristics and social resources needed for individuals and communities to access, understand, appraise, and use information and services to make decisions about health” [1]. From this, the National Health Service (NHS) of the UK emphasises two key factors for achieving better health literacy  [^1] https://www.england.nhs.uk/personalisedcare/health-literacy/, i.e., the individual’s comprehension ability and the health system itself. The “system” here refers to the complex network of health information and sources which promote it. These two factors are codependent. For instance, professionals write much healthcare information using complex language and terminologies without considering the readability of patients and the public in general. The health system must take into account the patient’s ability to achieve health literacy. Scientific studies have reported a correlation between low health literacy, poorer health outcomes, and poorer use of health care services [2, 3]. Thus, Plain Language Adaptation (PLA) of scientific reports in the healthcare domain is valuable for knowledge transformation and information sharing for public patients so as to promote public health literacy [4]. Nowadays, there have been industrial practices on such tasks, which include the publicly available plain summaries of scientific abstracts from the American College of Rheumatology (ACR) Virtual Meeting 2020 offered by the medicine company Novartis.com  [^2] https://www.novartis.com/node/65241.
The PLA task is related to text simplification and text summarisation, which are branches of the Natural Language Processing (NLP) field. This work investigates the biomedical domain PLA (BiomedPLA) using state-of-the-art large language models (LLMs) and control token methods [5, 6, 7, 8] that have been proven to be effective in such tasks. Examples of BiomedPLA can be seen in Figure 1 from the PLABA2023 shared task [^3] https://bionlp.nlm.nih.gov/plaba2023/. We highlight some of the factors in colours of such tasks, including sentence simplification in grey (removing clause “which” in the first sentence example; separating into two sentences and removing bracket for the second example), term simplification in yellow (“pharyngitis” and “pharynx” into “throat”), paraphrasing and synonyms in green (e.g. “acute” into “sore” and “posterior” into “back” for synonyms), and summarisation (on overall text in certain situations). The LLMs we applied include advanced Encoder-Decoder models (T5, SciFive, BART) and Generative Pre-trained Transformers (BioGPT, ChatGPT). The methodologies we applied include fine-tuning LLMs, prompt-based learning (PBL) on GPTs, and control token mechanisms on LLMs (BART-base and BART-large) with the efficient fine-tuning strategy. Using the publicly available PLABA (Plain Language Adaptation of Biomedical Abstracts) data set from [9], we demonstrate the capabilities of such models and carry out both quantitative and human evaluations of the model outputs. We also discuss the interesting findings from different evaluation metrics, their inconsistency, and future perspectives on this task. This work is a capstone based on our earlier findings reported in [10, 11].
Figure 1: Examples from the PLABA dataset on Biomedical Sentences Adaptation.
II Related Work
We first introduce recent developments in biomedical text simplification, then extend to broader biomedical LLMs related to this paper, followed by efficient training methodologies which we will apply to our work.
II-A Biomedical Text Simplification
To improve the health literacy level for the general public population, [12] developed the first lay language summarisation task using biomedical scientific reviews. The key points for this task include an explanation of context knowledge and expert language simplification. The evaluation included quality and readability using quantitative metrics. [13] carried out a survey, up to 2021, on biomedical text simplification methods and corpora using 45 relevant papers on this task, which data covers seven natural languages. In particular, the authors listed some published corpora on English and French languages and divided them into comparable, nonparallel, parallel, thesaurus, and pseudo-parallel. The quantitative evaluation metrics mentioned in these papers include SARI, BLEU, ROUGE, METEOR, and TER, among which three of them are borrowed from the machine translation (MT) field i.e., BLEU, METEOR, and TER [14]. Very recently, [15] transferred lay language style from expert physicians’ notes. They developed a comparable dataset from many non-parallel corpora on plain and expert texts. The baseline model applied for training is BART, with positive outcomes. [16] did a case study using chatGPT models on translating radiology reports into plain language for patient education. Detailed prompts were discussed to mitigate the GPT models to reduce “over-simplified” outputs and “neglected information”.
II-B Broader Biomedical LLMs
Beyond text simplification tasks, there have been active developments towards biomedical domain adaptation of LLMs in recent years. For instance, BioBERT used 4.5B words from PubMed and 13.5B words from PMC to do continuous learning based on the BERT pre-trained model. Then, it is fine-tuned in a task-specific setting on the following tasks: NER, RE, and QA [17]. In comparison, BioMedBERT [18] created new data sets called BREATHE using 6 million articles containing 4 billion words from different biomedical literature, mainly from NCBI, Nature Research, Springer Nature, and CORD-19, in addition to BioASQ, BioRxiv, medRxiv, BMJ, and arXiv. It reported better evaluation scores on QA data sets, including SQuAD and BioASQ, among other tested tasks, compared to other models.
BioMedLM 2.7B developed by Stanford Center for Research on Foundation Models (CRFM) and Generative AI company MOSAIC ML team  [^4] https://www.mosaicml.com/ formerly known as PubMedGPT 2.7B  [^5] https://huggingface.co/stanford-crfm/BioMedLM is trained on biomedical abstracts and papers using data from The Pile [19]. BioMedLM 2.7B claimed new state-of-the-art performance on the MedQA data set.
BioALBERT from [20] is based on the ALBERT [21] structure for training using biomedical data and reported higher evaluation scores on NER tasks of Disease, Drug, and Species, on several public data sets but much shorter training time in comparison to BioBERT. BioALBERT was also tested on broader BioNLP tasks using its different base and large models, including RE, Classification, Sentence Similarity, and QA by [22]
Afterwards, based on the T5 model structure [23], SciFive [24] was trained on PubMed Abstract and PubMed Central (PMC) data and claimed new state-of-the-art performance on biomedical NER and RE and superior results on BioAsq QA challenge over BERT and BioBERT. Similarly, BioBART [25] was developed recently based on the new learning structure BART model [26]. This work will examine T5, SciFive, and BART, leaving BioBART as a future work.
Other notable related works include a) the model comparisons in biomedical domains with different tasks by [27, 28, 29] on BERT, ALBERT, ELECTRA, PubMedBERT, and PubMedELECTRA; b) task- and domain-specific applications on QA by [30], on Medicines by [31], on radiology (RadBERT) by [32], concept normalisation by [33], abstract generation by [34]; c) language specific models such as in French [35] and Turkish [36]; and d) survey work by [37].
II-C Efficient Training
Due to the computational cost of the extra large-sized PLMs, some researchers proposed efficient training, which factor we will also apply to our study. These include some previously mentioned works.
[38] proposed Parameter-Efficient Transfer Learning for NLP tasks using their Adapter modules. In this method, the parameters of the original PLMs are fixed, and a few trainable parameters are added for each fine-tuning task, between 2-4 % of the original parameter sizes. Using GLUE benchmark data, they demonstrated that efficient tuning with the Adapter modules can achieve similar high performances compared to the BERT models with full fine-tuning of 100% parameters. ALBERT [21] applied parameter reduction training to improve the speed of BERT model learning. The applied technique uses a factorisation of the embedding parameters, which are decomposed into smaller-sized matrices before being projected into the hidden space. They also designed a self-supervised loss function to model the inner-sentence coherence. This reduced the parameter sizes from 108M in the BERT base to 12M in the ALBERT base models. Addressing similar issues, [39] proposed Prefix-tuning method, which modifies only 0.1% of the full parameters to achieve comparable performances using GPT-2 and BART for table-to-text generation and summarisation tasks.
Focus on the biomedical domain, [29] carried out fine-tuning stability investigation using the BLURB data set (Biomedical Language Understanding and Reasoning Benchmark) from [40]. Their findings show that freezing lower-level layers of parameters can be helpful for BERT-based model training, while re-initialising the top layers is helpful for low-resource text similarity tasks.
Instead of using Adapter modules [38] that require additional inference latency, [41] introduced Low-Rank Adaption (LoRA) that further reduces the size of trainable parameters by freezing the weights in PLMs and injects “trainable rank decomposition matrices” into every single layer of the Transformer structure for downstream tasks. The experiments were carried out on RoBERTa, DeBERTa, and GPTs that showed similar performances compared to the Adapter modules. We will apply LoRA for efficient fine-tuning on T5 and BioGPT in our work.
III Methodologies
@@
The overall framework of our experimental design is displayed in Figure 2. In the first step, we fine-tune selected LLMs including T5, SciFive, BioGPT, and BART, apply prompt-based learning for ChatGPTs, and optimise control mechanisms on the BART model. Then, we select the best performing two models using quantitative evaluation metrics SARI, BERTScore, BLEU and ROUGE. Finally, we chose a subset of the testing results of the two best-performing models for human evaluation.
@@
Figure 2: Model Development and Evaluation Pipeline. BART is fine-tuned using Wikilarge data. MAX step chooses the two best-performing models according to the automatic evaluation results using SARI and BERTScore.
III-A Models
@@
The models we investigated in our work include T5, SciFive, GPTs, BioGPT, BART, and Control Mechanisms; we will give more details below.
III-A1 T5
@@
T5 [23] used the same Transformer structure from [42] but framed the text-to-text learning tasks using the same vocabulary, sentence piece tokenisation, training, loss, and decoding. The pre-fixed tasks include summarisation, question answering, classification, and translation. The authors used the common crawl corpus and filtered to keep only natural text and de-duplication processing. They extracted 750GB of clean English data to feed into the model for multi-task pre-training. Different masking strategies are integrated into the T5 model to facilitate better performances of specific fine-tuning tasks. It has demonstrated state-of-the-art results across a wide spectrum of natural language processing tasks, showcasing its remarkable capabilities in capturing nuanced semantics and generating simplified texts while upholding high levels of accuracy. Notably, it has been successfully employed in various fields such as Clinical T5 by [43] and [44].
@@
In this paper, we fine-tuned three versions of T5, namely t5-small, t5-base, and t5-large, paired with their sentence-piece pre-trained tokenizer. Each is fine-tuned independently on the same dataset as the other models to provide comparable results. Note that we use the prompt “summarize:” as it is the closest to our task.
III-A2 SciFive
@@
Using the framework of T5, SciFive is a Large Language Model pre-trained on the biomedical domain and has demonstrated advanced performances on multiple biomedical NLP tasks [24].
@@
Similarly to our work on T5, we fine-tuned two versions of SciFive, namely SciFive-base and SciFive-large, paired with their pre-trained tokenizer. Each is fine-tuned independently on the same dataset as the other models to provide comparable results. We again use the prompt “summarize:” for task-relation purposes.
III-A3 OpenAI’s GPTs
@@
Given the remarkable performance demonstrated by OpenAI’s GPT models in text simplification[45], we decided to apply simplifications using “GPT-3.5-turbo” and GPT-4 via its API [^6] https://openai.com/blog/openai-api. Example prompts we used can be found in Figure 9.
III-A4 BioGPT
@@
BioGPT [46] is an advanced language model specifically designed for medical text generation. BioGPT is built upon GPT-3 but is specifically trained to understand medical language, terminology, and concepts. BioGPT follows the Transformer language model backbone and is pre-trained on 15 million PubMed abstracts. It has demonstrated a high level of accuracy and has great potential for applications in medicine. BioGPT is fine-tuned on the training and validation set, as with other encoder-decoder models (Figure 2).
III-A5 BART
@@
Like the default transformer structure, BART [26] aims to address the issues in BERT and GPT models by integrating their structures with a bi-directional encoder and an autoregressive decoder. In addition to the single token masking strategy applied in BERT, BART provides various masking strategies, including deletion, span masking, permutation, and rotation of sentences. Compared to GPT models, BART provides both leftward and rightward context in the encoders.
III-A6 Controllable Mechanisms
@@
We applied the modified control token strategy in [8] for both BART-base and BART-large models. The training includes 2 stages, leveraging both Wikilarge training set [47] and our split of training set from PLABA [48].
@@
They represent 1) the syntactic complexity, 2) the lexical complexity, 3) the inverse similarity of input and output at the letter level, and 4) the length ratio of input and output respectively.
@@
After fine-tuning, the next step is to find the optimal value of CTs. Following a similar process in MUSS [49], we applied Nevergrad [50] on the validation set to find the static optimal discrete value for DTD, WR, and LV. As for the LR, we applied the control token predictor to maximise the performance with the flexible value. The predictor is also trained on Wikilarge [47] to predict the potential optimal value for LR.
III-B LoRA and LLMs
@@
To evaluate bigger model architectures, we fine-tune FLAN-T5 XL [51] and BioGPT-Large, which have 3 billion and 1.5 billion parameters, respectively. FLAN-T5 XL is based on the pre-trained T5 model with instructions for better zero-shot and few-shot performance. To optimise training efficiency, and as our computational resources do not allow us to fine-tune the full version of these models, we employ the LoRA [41] technique, which allows us to freeze certain parameters, resulting in more efficient fine-tuning with minimal trade-offs.
III-C Metrics
We decide to evaluate our models using four quantitative metrics, namely BLEU [52], ROUGE [53], SARI [54], and BERTScore [55], each offering unique insights into text quality. SARI and BERTScore are used from EASSE [56] package; BLEU and ROUGE metrics are imported from the Hugging Face [^7] https://github.com/huggingface/evaluate implementations.
While BLEU quantifies precision by assessing the overlap between n-grams in the generated text and references, ROUGE measures recall by determining how many correct n-grams in the references are present in the generated text. This combination makes them useful as an initial indicative evaluation for machine translation and summarisation quality.
In contrast, SARI goes beyond n-gram comparisons and evaluates fluency and adequacy in translations. It does this by considering precision (alignment with references), recall (coverage of references), and the ratio of output length to reference length. SARI’s comprehensive approach extends its utility to broader evaluations of translation quality.
Finally, BERTScore delves into the semantic and contextual aspects of text quality. Using a pre-trained BERT model, it measures the similarity between word embeddings in the generated and reference texts. This metric provides insight into the semantic similarity and contextual understanding between generated and reference texts, making it akin to human evaluation. This metric does not quantify how good the simplification is but rather how much the meaning is preserved after simplification.
This comprehensive evaluation effectively addresses surface-level and semantic dimensions, resulting in a well-rounded and thorough assessment of the quality of machine-generated simplifications.
IV Experiments and Evaluations
PLABA data set is extracted from PubMed search results using 75 healthcare-related questions that MedlinePlus users asked. It includes 750 biomedical article Abstracts manually simplified into 921 adaptations with 7,643 sentence pairs in total. The dataset is publicly available via Zenodo  [^8] https://zenodo.org/record/7429310.
IV-A Data Preprocessing and Setup
@@
To investigate the selected models for training and fine-tuning, we divided the PLABA data into Train, Validation, and Test sets, aiming for an 8:1:1 ratio. However, in the final implementations, we found that there are only a few 1-to-0 sentence pairs, which might cause a negative effect in training the simplification models. Thus we eliminated all 1-to-0 sentence pairs. In addition, to better leverage the SARI score, we picked sentences with multi-references for validation and testing purposes. As a result, we ended up with the following sentence pair numbers according to the source sentences (5757, 814, 814).
IV-B Automatic Evaluation Scores